# MaxPoolBERT: Enhancing BERT Classification via Layer- and Token-Wise Aggregation

### v1

@@ Abstract
The [CLS] token in BERT is commonly used as a fixed-length representation for classification tasks, yet prior work has shown that both other tokens and intermediate layers encode valuable contextual information. In this work, we propose MaxPoolBERT, a lightweight extension to BERT that refines the [CLS] representation by aggregating information across layers and tokens. Specifically, we explore three modifications: (i) max-pooling the [CLS] token across multiple layers, (ii) enabling the [CLS] token to attend over the entire final layer using an additional multi-head attention (MHA) layer, and (iii) combining max-pooling across the full sequence with MHA. Our approach enhances BERT’s classification accuracy (especially on low-resource tasks) without requiring pre-training or significantly increasing model size. Experiments on the GLUE benchmark show that MaxPoolBERT consistently achieves a better performance on the standard BERT-base model.
@@ 1 Introduction
BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), is one of the best known Transformer-based (Vaswani et al., 2017) large language models. The core principle of BERT is the unsupervised pre-training approach on large corpora, enabling it to learn contextual word representations, which can then be used to solve various downstream tasks. Through fine-tuning, BERT adapts its representations to aggregate the most relevant information required for a given task.
A key component of BERT’s architecture is the classification token (abbreviated [CLS]), a special token that is prepended to every input sequence. During fine-tuning for downstream classification tasks, the [CLS] token serves as the only input to the classification head, which generates predictions for the task at hand. Through self-attention, the [CLS] token is expected to capture the sentence-level information necessary for downstream tasks. In this paper, we ask the question whether we can enrich the [CLS] with information from the layers below the top level.
Figure 1: MaxPoolBERT performs best on low-resource datasets. We show that MaxPoolBERT provides significant improvements for smaller datasets indicating that the model learns a better representation during fine-tuning.
Toshniwal et al. (2020) systematically compared different text span representations using BERT, and found that max-pooling performs quite well across tasks, though its effectiveness varies across tasks. Fine-tuning reduces the effect that different span representations had in the comparison (Stankevičius and Lukoševičius, 2024).
@@
Motivated by these findings, we explore incremental modifications to the BERT-base architecture for sequence classification, aiming to enhance its performance on downstream tasks. We specifically focus on improving the informativeness of the [CLS] token by (i) incorporating more width information of the whole sequence, and (ii) incorporating more depth information from additional layers. In the end we find that a mixture of these approaches leads to the best results.
@@
    We introduce MaxPoolBERT, a simple yet effective extension to BERT that enriches the [CLS] token representation using max-pooling and attention mechanisms across layers and tokens.
@@
    We show that our proposed approach improves fine-tuning performance on 7 out of 9 GLUE tasks and achieves an average gain of 1.25 points over the BERT-base baseline.
@@
Our best performing model will be made publicly available after the review process.
@@ 2 Related Work
Much research has been done dedicated to improving and optimizing BERT’s training process through architectural modifications and fine-tuning strategies. Below, we discuss advancements in fine-tuning stability, text representations, model enhancements, and training efficiency.
@@
The pre-training and fine-tuning paradigm for large language models such as BERT (Devlin et al., 2019) has led to significant improvements across a wide range of NLP tasks while keeping computational costs manageable. However, fine-tuning remains unstable due to challenges like vanishing gradients (Mosbach et al., 2021) and limited dataset sizes (Zhang et al., 2021). Several studies have proposed techniques to address this instability.
Zhang et al. (2021) explore re-initializing BERT layers before fine-tuning, demonstrating that retaining all pre-trained weights is not always beneficial for fine-tuning. They also show that extending fine-tuning beyond three epochs improves performance.
Hao et al. (2020) examine how fine-tuning affects BERT’s attention, finding that higher layers change significantly while lower layers remain stable. They propose a noise regularization method to enhance stability.
Mosbach et al. (2021) identify high learning rates as a key issue that cause fine-tuning instability. They propose using small learning rates with bias correction and increasing training iterations until nearly zero training loss is achieved.
Hua et al. (2021) introduce Layer-wise Noise Stability Regularization which further stabilizes fine-tuning through regularization.
Xu et al. (2023) propose self-ensemble and self-distillation mechanisms that enhance fine-tuning stability without requiring architectural changes or external data.
For an extensive, systematic review of methods for sentence-level embeddings from transformer models, we refer to Stankevičius and Lukoševičius (2024).
Faster Training.
In addition to stabilization, architectural enhancements have been introduced to boost BERT’s efficiency and effectiveness
Goyal et al. (2020) propose to eliminate tokens after fine-tuning, to reduce the time of inference. They found out, that in the highest layer of BERT-base, the token representations are very similar and carry similar information.
@@
While other approaches improve the input embedding size of BERT (Nussbaum et al., 2024) or refine the pre-training process for GPU’s (Geiping and Goldstein, 2023; Portes et al., 2023; Izsak et al., 2021), our work specifically concentrates on optimizing the [CLS] token during fine-tuning, leveraging the information captured in BERT’s layers after pre-training.
Improved BERT Fine-Tuning.
@@
Lastly there have been approaches to enhance the classification performance of BERT through optimized fine-tuning strategies and enriched sentence representations — areas that align closely with our approach, see also Stankevičius and Lukoševičius (2024) who provide a comprehensive survey of methods for extracting sentence-level embeddings from BERT.
Bao et al. (2021) construct sentence representations for classification by selecting meaningful n-grams and combining sub-tokens of a pre-trained BERT model into span representations using a max-pooling approach. These span-representations are max-pooled again into a sentence representation which is concatenated with the [CLS] token embedding.
Hu et al. (2024) develop a dynamically adjustable BERT architecture that optimizes both width (i.e., number of attention heads, number of neurons in the feed forward layer and sequence length) and depth (i.e., number of layers). Their method leverages knowledge distillation to maintain performance while reducing computational overhead.
Chang et al. (2023) introduce Multi-CLS BERT, a framework that fine-tunes BERT multiple times with varying random seeds and a multi-task objective to enhance performance.
Chen et al. (2023) present HybridBERT, which incorporates a hybrid pooling network and drop masking during pre-training to accelerate training and improve downstream accuracy. Their approach also includes local max pooling.
Sentence-BERT (Reimers and Gurevych, 2019) modifies the training objective to better capture semantic similarity at the sentence level, setting a benchmark for embedding-based classification.
Recently, Galal et al. (2024) compared multiple aggregation strategies for text classification, including an average pooling and a self-attention layer on top of the output embeddings. They also experiment with including the information of the last BERT layers into account. Their experiments are focused on Arabic classification tasks such as sentiment analysis on tweets. They find that freezing the BERT weights before fine-tuning boosts the classification results.
Similar to our experiments, Lehečka et al. (2020) experiment with a pooling operation on the whole sequence ot text to improve BERT for large-scale multi-label text classification. They find that combining max and mean pooling information leads to the greatest improvement.
3 Refining the [CLS] Token
BERT’s classification ([CLS]) token has a special function and serves as a summarizing token. In text classification tasks, it is the input to the classifier and is required to carry all the necessary information from the entire input sequence. This architecture is particularly efficient for fine-tuning and inference because only a single token is passed to the classifier.
However, it has been shown that other token representations in the layers of BERT also capture sentence-level representations (Rogers et al., 2020). We investigate whether the informativeness of the [CLS] token embedding can be further enhanced during fine-tuning, to improve downstream classification results. To do this, we include more depth information from other BERT layers and also more width information from other tokens within the sequence.
We study different versions of fine-tuning BERT for sequence classification tasks. All variants are described below.
3.1 Preliminaries.
Final-Layer [CLS] Representation.
As a baseline we use the [CLS] token of the final encoder layer of a pre-trained vanilla BERT-base model (Devlin et al., 2019) for classification (see Figure 2(a)). Recall that a single layer of BERT can be written as
$f_{i}:\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d},$ (1)
where $i$ indicates the layer number (for BERT-base we have 12 layers), $T$ is the number of tokens, and $d$ is the dimensionality of each token vector. We denote the values of the intermediate layers by $y^{(i)}$,
$y^{(1)}=f_{1}(x)\quad y^{(i+1)}=f_{i+1}(y^{(i)})$ (2)
[…]
The embedding of the [CLS] token serves as the input for the classification head $h$, which we choose to be just a linear layer without an activation function, since we are just interested in the plain expressiveness of the refinement (instead of adding tanh as in the original BERT implementation):
$h:\mathbb{R}^{1\times d}\rightarrow\mathbb{R}.$ (4)
$h\circ\text{CLS}\circ f_{12}\circ\cdots\circ f_{1}:\mathbb{R}^{T\times d}\rightarrow\mathbb{R}$ (5)
for BERT-base with 12 layers.
@@ 3 Refining the [CLS] Token
[…]
@@
Layer 1Layer NMulti-Head Attention LayerClassification HeadE[CLS]E[CLS]ETok1ETokTE[CLS]ETok1ETokT[CLS]Tok1Tok2TokT
@@
[…]
@@
Figure 2: Comparison of four BERT architectures for sequence classification. (Left above) Classical BERT for sequence classification architecture. (Right above) Maximizing the token embeddings of the [CLS] token over the last $k$ layers. (Left below) Adding an additional MHA layer before classification. (Right below) MaxPoolBERT architecture: The inputs are the same as in the standard BERT model. After the Nth layer (N = 12 for BERT-base), we apply a max-pooling operation over the last $k$ layers (we used $k=3$). The [CLS] token can then attend to every token after the max-pooling and the resulting [CLS] token embedding is used for classification.
Max pooling operation.
In order to utilize not only the last layer but several layers, we have to define a flexible maximizing operation, that can work with several sequences of vectors. For this, we write $\Theta_{t}^{(k)}\in\mathbb{R}^{k\times t\times d}$ for the tensor that contains the first $t$ token vectors (each $d$ dimensional) of the last $k$ layers. For instance, $\Theta_{1}^{(1)}\in\mathbb{R}^{1\times 1\times d}$ is the [CLS] token, and $\Theta_{1}^{(k)}\in\mathbb{R}^{k\times 1\times d}$ collects the token vectors the [CLS] token of the last $k$ layers. Similarly, $\Theta_{T}^{(1)}\in\mathbb{R}^{1\times T\times d}$ contains all token vectors of the last layer, and $\Theta_{T}^{(k)}\in\mathbb{R}^{k\times T\times d}$ all token vectors of the last $k$ layers.
Next, we define an element-wise max pooling operation that maximize over the first dimension, i.e.,
@@
As pytorch code, this is for $b$-sized minibatches of shape $b\times k\times t\times d$ the operation torch.max(Theta, dim=1).
3.2 Depth-Wise [CLS] Pooling ($\text{Max}_{\text{CLS}}$).
[…]
3.3 Token-Wise Attention via Additional MHA Layer.
An alternative way to enrich the information in the [CLS] token, is to consider horizontal information (i.e., more width, see Figure 2(c)). For this, we include all tokens of the last layer. To obtain a single vector, we employ an additional multi-head attention (MHA) layer on encoder output, but compute the attention only for the [CLS] token. Writing the MHA as usual (see (Vaswani et al., 2017)),
[…]
@@
we note that in our case $Q=\texttt{[CLS]}_{12}$ and $K=V=y^{(12)}$. Thus the [CLS] token can attend to all other tokens once more before classification. Note that the additional MHA layer is not part of the pre-training process and is only added and initialized before the fine-tuning process. We use the default initialization of the Pytorch [^1] https://pytorch.org/ multi-head attention implementation which is a Xavier uniform initialization (Glorot and Bengio, 2010). For the number of heads we choose $h=4$.
3.4 Sequence-Wide Pooling with MHA ($\text{Max}_{\text{Seq}}$ + MHA).
Finally, we combine the additional depth and width information of $\text{Max}_{\text{CLS}}$ and MHA by extending the max-pooling operation to whole sequences of the last $k$ layers by using $\max(\Theta_{t}^{(k)})\in\mathbb{R}^{k\times t\times d}$. We call this setup $\text{Max}_{\text{Seq}}$+ MHA, since the maximum is now along the whole sequence and the additional MHA layer aggregates the pooled information. We call this approach MaxPoolBERT in the following.
We did consider other variations of these ideas as well: for instance instead a component-wise maximum operation, we calculated the norms of the token vectors and used those to select among the layers. However, the performance was not significantly different to the reported results (see Appendix).
@@
Model CoLA MRPC QNLI MNLI SST-2 STS-B RTE QQP WNLI
MCC Acc. F1 Acc. m mm Acc. Sp. Acc. Acc. F1 Acc.
Train Size 8.5k 3.7k 3.7k 105k 393k 393k 67k 5.75k 2.5k 364k 364k 634
Baseline 53.59 82.43 87.49 90.96 84.27 84.57 92.55 88.47 63.42 90.65 87.40 49.77
MHA 55.88 83.38 89.51 90.49 84.37 84.67 92.32 88.04 64.98 90.67 87.45 55.4
$\text{Max}_{\text{Seq}}$+MHA 55.35 85.95 89.78 90.73 83.82 84.24 92.74 88.22 66.06 90.65 87.59 55.4
$\Delta$ 2.29 3.52 2.29 0.19 0.24 0.1 0.19 0.5 3.61 0.02 0.19 5.63
Table 2: Performance on GLUE validation tasks (average of 3 seeds). The size of the training data set is highlighted in gray. We report Matthews correlation coefficient (MCC) for CoLA, accuracies for matched (m) and mismatched results (mm) for MNLI, and Spearman correlation (Sp.) for STS-B. Below we report the improvement from one variant over the baseline as $\Delta$.
Model GLUE avg.
Baseline 79.63
$\text{Max}_{\text{CLS}}$ 80.02
MHA 80.70
$\text{Max}_{\text{Seq}}$+MHA 80.88
$\Delta$ 1.25
Table 3: Average performance across all GLUE tasks. MaxPoolBERT shows a consistent gain over BERT-base.

### v2

@@ Abstract
The [CLS] token in BERT is commonly used as a fixed-length representation for classification tasks, yet prior work has shown that both other tokens and intermediate layers encode valuable contextual information. In this work, we study lightweight extensions to BERT that refine the [CLS] representation by aggregating information across layers and tokens. Specifically, we explore three modifications: (i) max-pooling the [CLS] token across multiple layers, (ii) enabling the [CLS] token to attend over the entire final layer using an additional multi-head attention (MHA) layer, and (iii) combining max-pooling across the full sequence with MHA. Our approach, called MaxPoolBERT, enhances BERT’s classification accuracy (especially on low-resource tasks) without requiring new pre-training or significantly increasing model size. Experiments on the GLUE benchmark show that MaxPoolBERT consistently achieves a better performance than the standard BERT base model on low resource tasks of the GLUE benchmark.
@@ 1 Introduction
BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019), is one of the best known Transformer-based (Vaswani et al., 2017) language models. The core principle of BERT is the unsupervised pre-training approach on large corpora, enabling it to learn contextual word representations, which can then be used to solve various downstream tasks. Through fine-tuning, BERT adapts its representations to aggregate the most relevant information required for a given task.
Figure 1: MaxPoolBERT performs best on low-resource datasets. We show that our methods, in particular MaxPoolBERT, provide significant improvements for smaller datasets indicating that the model learns a better representation during fine-tuning (top-left).
A key component of BERT’s architecture is the classification token (abbreviated [CLS]), a special token that is prepended to every input sequence. During fine-tuning, the [CLS] token serves as the only input to the classification head, which generates predictions for the task at hand. Through self-attention, the [CLS] token is expected to capture the sentence-level information necessary for downstream tasks. In this paper, we ask the question whether we can enrich the [CLS] token with information from the layers below the top level.
@@
Motivated by these findings, we explore incremental modifications to the BERT base architecture for sequence classification, aiming to enhance its performance on downstream tasks. We specifically focus on improving the informativeness of the [CLS] token by (i) incorporating more width information of the whole sequence, and (ii) incorporating more depth information from additional layers. In the end we find that a mixture of these approaches leads to the best results.
@@
    We introduce MaxPoolBERT, an effective extension to BERT that enriches the [CLS] token representation using max-pooling and attention mechanisms across layers and tokens.
@@
    We show that our proposed approach improves fine-tuning performance on 7 out of 9 GLUE tasks and achieves an average gain of 1.25 points over the BERT base baseline.
@@
All of our models will be made publicly available after the review process.
@@ 2 Related Work
Much research has been done dedicated to improving and optimizing BERT’s training process through architectural modifications and fine-tuning strategies. Below, we discuss advancements in fine-tuning stability, text representations, model enhancements, and training efficiency. Our work falls within the branch of research aimed at optimizing BERT’s representation to enhance downstream classification results, with a particular focus on augmenting the informativeness of the [CLS] token.
@@
The pre-training and fine-tuning paradigm for language models such as BERT (Devlin et al., 2019) has led to significant improvements across a wide range of NLP tasks while keeping computational costs manageable. However, fine-tuning remains unstable due to challenges like vanishing gradients (Mosbach et al., 2021) and limited dataset sizes (Zhang et al., 2021). Several studies have proposed techniques to address this instability.
Zhang et al. (2021) explore re-initializing BERT layers before fine-tuning, demonstrating that retaining all pre-trained weights is not always beneficial for fine-tuning. They also show that extending fine-tuning beyond three epochs improves performance. Hao et al. (2020) examine how fine-tuning affects BERT’s attention, finding that higher layers change significantly while lower layers remain stable. They propose a noise regularization method to enhance stability. Mosbach et al. (2021) identify high learning rates as a key issue that cause fine-tuning instability. They propose using small learning rates with bias correction and increasing training iterations until nearly zero training loss is achieved. Hua et al. (2021) introduce Layer-wise Noise Stability Regularization which further stabilizes fine-tuning through regularization. Xu et al. (2023) propose self-ensemble and self-distillation mechanisms that enhance fine-tuning stability without requiring architectural changes or external data.
Our method, while not explicitly targeting stability, contributes to more robust performance especially on low-resource tasks by enabling the [CLS] token to integrate a broader context via pooling and attention. We analyze fine-tuning stability of our variants in Section 5.2.
Faster and More Efficient Training.
In addition to stabilization, architectural enhancements have been introduced to boost BERT’s efficiency and effectiveness. Goyal et al. (2020) propose to eliminate tokens after fine-tuning, to reduce the time of inference. They discovered that the token representations in the highest layer of BERT base carry similar information.
@@
While other approaches improve the input embedding size of BERT (Nussbaum et al., 2024) or refine the pre-training process for GPU’s (Geiping & Goldstein, 2023; Portes et al., 2023; Izsak et al., 2021), our work specifically concentrates on optimizing the [CLS] token during fine-tuning, leveraging the information captured in BERT’s layers after pre-training.
Improved BERT Fine-Tuning and Representation Learning.
@@
Lastly, several approaches refine BERT’s classification capability through optimized fine-tuning strategies and enriched sentence representations - areas that align closely with our approach (see also Stankevičius & Lukoševičius (2024) who provide a comprehensive survey of methods for extracting sentence-level embeddings from BERT).
Toshniwal et al. (2020) systematically compared different text span representations using BERT, and found that max-pooling performs quite well across tasks, though its effectiveness varies. Bao et al. (2021) construct sentence representations for classification by selecting meaningful n-grams and combining sub-tokens of a pre-trained BERT model into span representations using a max-pooling approach. In contrast, our method does not require span selection or input modification, and applies pooling and attention directly to hidden states during fine-tuning.Hu et al. (2024) introduce a flexible BERT architecture with dynamic width and depth that adapts the number of attention heads, hidden size, and number of layers at inference time using knowledge distillation. Our approach does not alter the base architecture, instead we enrich the fixed-size [CLS] embedding to boost classification performance.
Chang et al. (2023) introduce Multi-CLS BERT, a framework that modifies pre-training and adds multiple [CLS] tokens to the sequence for fine-tuning. We achieve comparable results on the GLUE benchmark without altering the pre-training setup. Chen et al. (2023) present HybridBERT, which incorporates a hybrid pooling network and drop masking during pre-training to accelerate training and improve downstream accuracy. While HybridBERT combines multiple pooling strategies (mean, max, and attention) to replace the [CLS] token, we retain the original [CLS] embedding and instead enrich it through architectural refinements such as an additional multi-head attention layer and optional sequence-wide pooling. This allows our method to be applied to any pre-trained BERT-like model without re-training, with particular benefits observed on tasks with limited training data.
Recently, Galal et al. (2024) explore aggregation techniques such as mean pooling and self-attention on output embeddings for Arabic sentiment analysis. They show that freezing BERT during fine-tuning can boost performance. Our method can be combined with such techniques but focuses on improving the [CLS] pathway, especially under low-resource conditions.
Lastly, Lehečka et al. (2020) propose modifying BERT’s output pooling strategy to improve large-scale multi-label text classification. Specifically, they replace the [CLS] token with combined mean and max pooling over the final hidden states of all tokens, arguing that this captures richer semantic information for classification. While their method entirely discards the [CLS] embedding, our approach retains it and enhances its contextual richness by integrating sequence-wide information via additional architectural layers during fine-tuning.
@@
Layer 1Layer NMaxClassification HeadE[CLS]E[CLS]ETok1ETokTE[CLS]ETok1ETokT[CLS]Tok1Tok2TokTkk
@@
Layer 1Layer NMHA LayerClassification HeadE[CLS]E[CLS]ETok1ETokTE[CLS]ETok1ETokT[CLS]Tok1Tok2TokT
@@
Layer 1Layer NMaxMHA LayerClassification HeadE[CLS]E[CLS]ETok1ETokTE[CLS]ETok1ETokT[CLS]Tok1Tok2TokTkk
@@
Figure 2: Comparison of four BERT architectures for sequence classification. (Left above) Classical BERT for sequence classification architecture. (Right above) Applying max-pooling on the token embeddings of the [CLS] token over the last $k$ layers. (Left below) Adding an additional MHA layer before classification. (Right below) MaxPoolBERT architecture: After the Nth layer (N = 12 for BERT base), we apply a sequence-wide max-pooling operation over the last $k$ layers (we used $k=3$). The [CLS] token can then attend to every token after the max-pooling and the resulting [CLS] token embedding is used for classification.
3 Refining the [CLS] Token
It has been shown that other token representations in the layers of BERT also capture sentence-level representations (Rogers et al., 2020). We investigate whether the informativeness of the [CLS] token embedding can be further enhanced during fine-tuning, to improve downstream classification results. To do this, we include more depth information from other BERT layers and also more width information from other tokens within the sequence. We study different versions of fine-tuning BERT for sequence classification tasks. All variants are described below.
3.1 Preliminaries
Final-Layer [CLS] Representation.
As a baseline we use the [CLS] token of the final encoder layer of a fine-tuned vanilla BERT base model (Devlin et al., 2019) for classification (see Figure 2(a)). Recall that a single layer of BERT can be written as
$f_{i}:\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d},$ (1)
where $i$ indicates the layer number (BERT base has 12 layers), $T$ is the number of tokens, and $d$ is the dimensionality of each token vector. We denote the values of the intermediate layers by $y^{(i)}$:
$y^{(1)}=f_{1}(x),\quad\quad y^{(i+1)}=f_{i+1}(y^{(i)}).$ (2)
$\texttt{[CLS]}_{i}=t_{1i}\in\mathbb{R}^{1\times d}.$ (3)
The embedding of the [CLS] token serves as the input for the classification head $c$, which we have choosen to be a simple linear layer without an activation function, since we are just interested in the plain expressiveness of the refinement (instead of adding tanh as in the original BERT implementation):
$c:\mathbb{R}^{1\times d}\rightarrow\mathbb{R}.$ (4)
$(c\circ\text{CLS}\circ f_{12}\circ\cdots\circ f_{1}):\mathbb{R}^{T\times d}\rightarrow\mathbb{R}$ (5)
for BERT base with 12 layers.
Max-pooling operation.
The final layers of a BERT model are known to contain the task-specific information. In order to utilize not only the last layer but several layers, we have to define a flexible maximizing operation, that can work with several sequences of vectors. For this, we write $\Theta_{t}^{(k)}\in\mathbb{R}^{k\times t\times d}$ for the tensor that contains the first $t$ token vectors (each $d$ dimensional) of the last $k$ layers. For instance, $\Theta_{1}^{(1)}\in\mathbb{R}^{1\times 1\times d}$ is the [CLS] token, and $\Theta_{1}^{(k)}\in\mathbb{R}^{k\times 1\times d}$ collects the token vectors the [CLS] token of the last $k$ layers. Similarly, $\Theta_{T}^{(1)}\in\mathbb{R}^{1\times T\times d}$ contains all token vectors of the last layer, and $\Theta_{T}^{(k)}\in\mathbb{R}^{k\times T\times d}$ all token vectors of the last $k$ layers.
Next, we define an element-wise max-pooling operation that maximizes over the first dimension, i.e.,
@@ 3.1 Preliminaries
Written as Pytorch [^1] https://pytorch.org/ code, the operation is torch.max(Theta, dim=1) for $b$-sized minibatches of shape $b\times k\times t\times d$.
Mean-pooling operation.
Several studies indicate, that max-pooling seems to be a stable choice to aggregate information into a single sentence representation. In the experimental section (Section 4), we also consider mean-pooling to challenge these results. For this, we apply an element-wise mean-pooling operation
on every vector of our $k$ chosen layers (defined analogously as the max-pooling operation).
3.2 Depth-Wise [CLS] Pooling ($\text{Max}_{\text{CLS}}$)
To use the vertical information (i.e., more depth) as one possible improvement for BERT’s fine-tuning, we take information from the last $k$ layers (instead of only from the last layer): we extract the last $k$ [CLS] embeddings $\texttt{[CLS]}_{12-k+1},\dots,\texttt{[CLS]}_{12}]$ which corresponds to $\Theta_{1}^{(k)}\in\mathbb{R}^{k\times 1\times d}$ (using the notation of the previous paragraph). Then we apply the element-wise max-pooling operation on the extracted tokens (see Figure 2(b)).
3.3 Token-Wise Attention via Additional MHA Layer (MHA)
The orthogonal way to enrich the information in the [CLS] token, is to consider horizontal information (i.e., more width, see Figure 2(c)). For this, we include all tokens of the last layer. To obtain a single vector, we employ an additional multi-head attention (MHA) layer on the encoder output, but compute the attention only for the [CLS] token. We write the MHA as (see Vaswani et al., 2017),
$\text{MHA(Q,K,V)}=[\text{head}_{1},\dots,\text{head}_{h}]W_{0}$ (8)
@@ 3.3 Token-Wise Attention via Additional MHA Layer (MHA)
Using the standard BERT base model with 12 layers, we have $Q=\texttt{[CLS]}_{12}$ and $K=V=y^{(12)}$. Through the attention mechanism, the [CLS] token can attend to all other tokens once more before classification. Note that the additional MHA layer is not part of the pre-training process and is added and initialized before the fine-tuning process. We use the default initialization of the Pytorch1 multi-head attention implementation which is a Xavier uniform initialization (Glorot & Bengio, 2010). For the number of attention heads we choose $h=4$.
@@
3.4 Sequence-Wide Pooling with MHA ($\text{Max}_{\text{Seq}}$ + MHA & $\text{Mean}_{\text{Seq}}$ + MHA )
Finally, we combine the additional depth and width information of $\text{Max}_{\text{CLS}}$ and MHA by extending the max-pooling operation to the whole sequences of the last $k$ layers by using $\max(\Theta_{T}^{(k)})\in\mathbb{R}^{k\times T\times d}$. We call this setup $\text{Max}_{\text{Seq}}$+ MHA, since the maximum is now along the whole sequence and the additional MHA layer aggregates the pooled information. We call this approach MaxPoolBERT in the following. As a variant, we replaced max pooling with mean pooling. We report the results for mean pooling with an additional MHA layer as $\text{Mean}_{\text{Seq}}$+ MHA.